The number on two consecutively numbered gym lockers have a sum of 131. What are the locker numbers?
Found 2 solutions by Alan3354, macston:
Answer by Alan3354(69443)

131/2 = 65.5
--> 65 & 66

Answer by macston(5194)

.
x and x+1 are consecutive locker numbers.
x+(x+1)=131
2x+1=131
2x=130
x=65 ANSWER 1: The lower number is 65.
x+1=65+1=66 ANSWER 2: The higher number is 66.